OpenStudy (anonymous):

(all are log base 5) log64-log8/3+log2=log4p solve for p round to 4 decimal places if ness.

OpenStudy (turingtest):

\[\log_5(64)-\log_5({8\over3})+\log_5(2)=\log_5(4p) \]right?

OpenStudy (anonymous):

correct

OpenStudy (turingtest):

\[=\log_5(64*{3\over8}*2)=\log_5(4p)\to4p=48\to p=12\]

OpenStudy (turingtest):

I just used the rules loga-logb=log(a/b) and loga+logb=log(ab) then raised 5 to each side

OpenStudy (anonymous):

okay thanks this really helps.. the mistake i made was instead of multiplying (64*3/8*2) i did this (64-3/8+2) im not really great at math

OpenStudy (turingtest):

You just gotta learn the log rules: \[\log a = \log b=\log(ab)\]\[\log(a^n)=n \log a\]

OpenStudy (anonymous):

notice that the base is entirely irrelevant to this question

OpenStudy (anonymous):

are the bases only irrelevant to the problem if theyre all the same? and @turintest thats one thing they havent told me in my online class... thanks!

OpenStudy (anonymous):

@turingtest. shouldnt it be log a + log b = Log (AB)?

OpenStudy (turingtest):

yes you are right. just making sure you noticed ;)

OpenStudy (anonymous):

\[\log_{2}N= 1/4\log_{2}16+ 1/2 \log_{2} 49 \]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

do i start with putting the exponents back or leave it as is?

OpenStudy (turingtest):

Notice we can use \[n \log a=\log(a^n)\]to get closer to having everything under one log sign, so try that first...

OpenStudy (turingtest):

yes that is "putting the exponents back" as you call it.

OpenStudy (anonymous):

\[\log_{2}16^{1/4}+ \log_{2} 49^{1/2}\]

OpenStudy (anonymous):

then i would take the natural log of each?

OpenStudy (turingtest):

Whoa, let's just look at the exponential fractions first: \[16^{1/4}\] What does that equal? Hint: Remember that\[ x^{a/b}=\sqrt[b]{x^a}\]

OpenStudy (anonymous):

\[\sqrt[4]{16^1} \]=1.091

OpenStudy (turingtest):

No you used a calculator improperly or something. Try to apply that rule above to \[49^{1/2}\] first, it's easier.

OpenStudy (anonymous):

7?

OpenStudy (turingtest):

Exactly, it's a square root in disguise. That means the other is a fourth root. What is the fourth root of 16? (it's a whole number)

OpenStudy (anonymous):

2

OpenStudy (turingtest):

Right, so our expression becomes...

OpenStudy (anonymous):

\[\log_{2}N= \log_{2}2+\log_{2}7 \]

OpenStudy (turingtest):

perfect, now apply our other rule \[\log a + \log b = \log (ab)\]

OpenStudy (anonymous):

\[\log_{2}N= \log_{2}14 \]

OpenStudy (turingtest):

Awesome, almost done! We just have to realize that, once again, because the bases are the same they don't matter! When we raise 2 to the power of both sides the logs cancel, so N=?

OpenStudy (anonymous):

n=14

OpenStudy (turingtest):

You got it, congrads, I've gotta have lunch ;)

OpenStudy (anonymous):

lol feel free to eat! thanks so much for helping me and explaining everything. your the most help ive gotten with this class in 2 years (long story) but thanks again! its much appreciated

OpenStudy (turingtest):

no prob happy to help!

OpenStudy (anonymous):

:) thanks again!